# A Review of LLM-Assisted Ideation

### v2

@@ 4.4 The Iteration of LLM-assisted Ideation Stages
Three studies [huang2023causalmapper_G5, gonzalez2024collaborative_G7, brade2023promptify_M3] specifically involve selecting and processing individual ideas from idea clusters or multiple ideas. In these cases, the iterative process returns from the multi-idea evaluation and selection stage to the individual idea evaluation stage. This indicates that filtering and evaluating promising ideas at the idea cluster level is a promising approach.
Interestingly, two studies [pu2024ideasynth_S1, liu2024personaflow_S2] focus on the ideation of scientific research ideas. In these cases, idea iteration involves the exploration of new ideas and research starting points from multiple research ideas, representing a return to the scope definition stage of the ideation process. This finding suggests that in the context of scientific research, iterative ideation may involve revisiting and redefining the scope of the problem or research question, allowing for the exploration of novel research directions.
One study [kocaballi2023conversational_D1] employs the creation of different roles to facilitate product ideation. This approach represents an iteration that returns from the idea evaluation stage to the foundational materials collection stage, where the fundamental elements and perspectives that inform the ideation process are established. By iterating on the roles and personas involved in the ideation activity, this study highlights the importance of considering diverse viewpoints and contexts in the development of innovative product ideas.
@@ 5.1 Overview of Interaction Design in LLM-assisted Ideation Systems
[heyman2024supermind_G2] tool G2 Individual Text board & sidebar - Triggered Functions
@@
[xu2024jamplate_G3] tool G3 Individual Text board Add & Edit User-defined Prompt & Triggered Functions
[di2022idea_G4] tool G4 Individual Text board - Triggered Functions
@@
[huang2023causalmapper_G5] tool G5 Individual Text Canvas Add & Delete & Edit Triggered Functions
@@
[gonzalez2024collaborative_G7] tool G7 Group Text Canvas & Chatbot Add & Delete & Edit Open Interaction & User-defined Prompt
@@
[aikawa2023introducing_G8] tool G8 Group Text & Audio & Physicalisation Augmented Reality - Triggered Functions
[chen2024asknaturenet_G10] tool G10 Individual Text Canvas - Triggered Functions
@@
[suh2024luminate_G13] tool G13 Individual Text Text Editor & Canvas Add & Edit User-defined Prompt & Triggered Functions
@@
[kang2024biospark_G14] tool G14 Individual Text & Image board & sidebar - Triggered Functions
[blazevic2024real_G15] tool G15 Group Text board - Triggered Functions
@@
[lee2024generating_D7] tool D7 Individual Text & Image Config Panel & Chatbot - Open Interaction
@@
[lu2024large_D9] tool D9 Individual Text & Animation board & canvas - Automated Workflows
@@
[hou2024c2ideas_D13] tool D13 Individual Text & Image Chatbot & Config Panel & Canvas Edit Open Interaction & User-defined Prompt
@@
[huang2024plantography_D14] tool D14 Individual Text & Image Config Panel & Canvas Add & Delete & Edit Triggered Functions
@@
[goodman2022lampost_W1] tool W1 Individual Text Text Editor & Sidebar Add & Delete & Edit User-defined Prompt & Triggered Functions
@@
[kim2024towards_W2] tool W2 Individual Text Text Editor & Sidebar & Chatbot - Open Interaction & User-defined Prompt & Triggered Functions
@@
[chung2022talebrush_W3] tool W3 Individual Text & Sketch Text Editor & Canvas Add & Delete Triggered Functions
@@
[zhang2023visar_W4] tool W4 Individual Text Text Editor & Canvas Add & Delete & Edit User-defined Prompt & Triggered Functions
@@
[yuan2022wordcraft_W5] tool W5 Individual Text Text Editor & Sidebar Add & Delete & Edit User-defined Prompt & Triggered Functions
[petridis2023anglekindling_W6] tool W6 Individual Text board & sidebar - Predefined Prompt
[ghajargar2022redhead_W7] tool W7 Individual Text Text Editor & Canvas - Triggered Functions
[wan2024felt_W8] tool W8 Individual Text Chatbot - Open Interaction
@@
[kim2023cells_W9] tool W9 Individual Text Config Panel & Text Editor Edit User-defined Prompt & Triggered Functions
[goldi2024intelligent_W10] tool W10 Individual Text Text Editor & Sidebar - Triggered Functions
@@
[reza2024abscribe_W13] tool W13 Individual Text Text Editor & Sidebar Add & Delete & Edit User-defined Prompt & Triggered Functions
@@
[benharrak2024writer_W14] tool W14 Individual Text Text Editor & Sidebar Add & Delete & Edit User-defined Prompt & Triggered Functions
@@
[pu2024ideasynth_S1] tool S1 Individual Text board & Canvas Add & Delete & Edit User-defined Prompt & Triggered Functions
@@
[liu2024personaflow_S2] tool S2 Individual Text Canvas Add & Delete & Edit User-defined Prompt & Triggered Functions
@@
[radensky2024scideator_S3] tool S3 Individual Text board Add & Delete User-defined Prompt & Triggered Functions
[cai2023designaid_M1] tool M1 Individual Text & Image board - Triggered Functions
[wang2023popblends_M2] tool M2 Individual Text & Image board - Predefined Prompt
@@
[brade2023promptify_M3] tool M3 Individual Text & Image Config Panel & Canvas - User-defined Prompt & Triggered Functions
@@
[wang2024lave_M4] tool M4 Individual Text & Video Canvas & Chatbot - Open Interaction & User-defined Prompt & Triggered Functions
@@
[qian2024shape_M5] tool M5 Individual Text & Animation Config Panel & Canvas Edit Automated Workflows
[rosenberg2023conversational_G6] framework G6 Group Text & Audio - - Automated Workflows
[besta2024graph_G12] framework G12 Individual Text - - Automated Workflows
[bhavya2023cam_G16] framework G16 Individual Text - - Open Interaction & Predefined Prompt
[dharuman2023protein_D5] framework D5 Individual Text - - Automated Workflows
[lanzi2023chatgpt_D6] framework D6 Group Text - - Automated Workflows
[lee2024gpt_S6] framework S6 Individual Text - - Automated Workflows
[chen2024foundation_M8] framework M8 Individual Text & Image - - Open Interaction
[shaer2024ai_G1] method G1 Group Text - - Open Interaction
[schuller2024generating_G9] method G9 Individual Text - - Open Interaction
[bernstein2024like_G11] method G11 Individual Text - - Open Interaction
[paredes2024creative_G17] method G17 Individual Audio & Physicalisation - - Open Interaction
[lee2024prompt_G18] method G18 Individual Text - - Open Interaction
[kocaballi2023conversational_D1] method D1 Individual Text - - Open Interaction
[bilgram2023accelerating_D2] method D2 Group Text - - Open Interaction
[sudhakaran2023prompt_D3] method D3 Individual Text - - Open Interaction
[todd2023level_D4] method D4 Individual Text - - Open Interaction
[wang2023task_D8] method D8 Individual Text - - Open Interaction & Predefined Prompt
[ege2024chatgpt_D10] method D10 Group Text - - Open Interaction
[junior2023chatgpt_D11] method D11 Group Text - - Open Interaction
[veloso2024forming_D12] method D12 Individual Text & Image - - Open Interaction
[li2024mystery_W11] method W11 Individual Text - - Open Interaction
[pividori2024publishing_W12] method W12 Individual Text - - Open Interaction
[yuan2024llmcrit_W15] method W15 Individual Text - - Open Interaction
[banker2024machine_S4] method S4 Individual Text - - Open Interaction
[chen2024use_S5] method S5 Individual Text - - Open Interaction
[…]
[wu2023iconshop_M7] method M7 Individual Text & Image - - Open Interaction
@@ 5.3 User Group
The prevalence of individually-oriented systems, such as LaMPost [goodman2022lampost_W1] and LAVE [wang2024lave_M4], demonstrates the emphasis on leveraging LLMs to enhance personal creativity and idea generation. These systems provide users with powerful tools to independently generate, refine, and evaluate ideas, augmenting their creative capabilities and enabling them to explore a wider range of possibilities.
However, the importance of group ideation cannot be overlooked, particularly in the context of large enterprises and organizations. Collaborative ideation offers several key benefits, including the exchange of diverse perspectives, the fostering of constructive criticism, and the cultivation of a shared sense of ownership over the generated ideas [gonzalez2024collaborative_G7]. By bringing together individuals with different backgrounds, expertise, and viewpoints, team-based ideation can lead to more robust and innovative solutions.
Despite the critical role of group ideation, the interaction table reveals a notable scarcity of LLM-assisted tools specifically designed to support collaborative processes. Among the reviewed papers, only a few systems, such as the Collaborative Canvas [gonzalez2024collaborative_G7] and the Conversational Swarm Intelligence system [rosenberg2023conversational_G6], are capable of supporting the collaborative processes in group ideation. This also indicates that developing collaborative systems for group ideation still faces unique challenges, which may include developing tools that support synchronous and asynchronous ideation, managing social dynamics, facilitating turn-taking, integrating the contributions of multiple participants, and ensuring equitable participation. There exist significant opportunities for future research and development. By focusing on the design and implementation of systems that cater to the specific needs of collaborative ideation, researchers can fully harness the potential of LLMs in supporting creative processes within teams and organizations.
@@ 5.4 Interaction Modality
The prevalence of text-based interaction in LLM-assisted ideation systems can be attributed to the initial development of LLMs, which focused heavily on text-based applications. This has led to the extensive use of LLMs in domains such as writing and other text-centric tasks. Many of the reviewed papers, such as those dealing with argumentative writing [zhang2023visar_W4] and conceptual involvement [wang2023task_D8], rely solely on text-based interaction to facilitate ideation processes.
However, in the realm of multimedia creation or domains involving multimedia elements, there is a growing trend towards combining text with various forms of media for interaction. For instance, image generation tasks [brade2023promptify_M3] often employ a combination of text and image-based interaction, while video editing tasks [wang2024lave_M4] incorporate text and video modalities. Group ideation meetings [rosenberg2023conversational_G6] may leverage text and audio interaction to facilitate collaboration. Other interactive modalities explored in the reviewed papers include text and sketch [chung2022talebrush_W3], text and animation [lu2024large_D9, qian2024shape_M5], and even text and bodily movements [aikawa2023introducing_G8].
Interestingly, only one paper in the reviewed set utilizes a non-text-based interaction modality. [paredes2024creative_G17] explores the concept of assisting idea incubation in the context of commuter driving, employing a combination of driver body state and audio interaction. This unique approach highlights the potential for LLM-assisted ideation systems to extend beyond traditional text-based interactions and adapt to specific contextual requirements.
@@ 5.6 Idea Management Capability
Our analysis revealed that 17 tools (50%) developed idea components that considered users’ ability to manage ideas. Among these 17 tools, 10 incorporated all three functional components of adding, deleting, and editing ideas [huang2023causalmapper_G5, huang2024plantography_D14, reza2024abscribe_W13]. By offering a comprehensive set of idea management capabilities, these tools empower users to fully engage with the generated ideas, refining and shaping them to align with their creative goals.
Two tools [xu2024jamplate_G3, suh2024luminate_G13] focused on the addition and modification of ideas, while another two tools [chung2022talebrush_W3, radensky2024scideator_S3] prioritized the addition and deletion of ideas. These variations in idea management capabilities reflect the different emphases and objectives of each tool, catering to specific aspects of the ideation process.
Interestingly, three tools [hou2024c2ideas_D13, kim2023cells_W9, qian2024shape_M5] solely considered the modification of ideas. While the ability to edit ideas is undoubtedly valuable, the absence of addition and deletion functionalities may limit users’ capacity to fully explore the creative space and manage the generated ideas effectively.
@@ 5.7 Human-LLM Interaction Format
Predefined prompts are used in 4 out of 61 reviewed papers (7%). Two tool-type studies [petridis2023anglekindling_W6, wang2023popblends_M2] incorporate predefined prompts into their backend systems. The other two studies [bhavya2023cam_G16, wang2023task_D8] involve users directly interacting with LLMs using structured, predefined prompts.
@@ 6.1 User Study Design
One Condition - - [shaer2024ai_G1], [xu2024jamplate_G3], [rosenberg2023conversational_G6], [bernstein2024like_G11], [suh2024luminate_G13], [kang2024biospark_G14], [blazevic2024real_G15], [bhavya2023cam_G16], [kocaballi2023conversational_D1], [lanzi2023chatgpt_D6], [hou2024c2ideas_D13], [kim2024towards_W2], [chung2022talebrush_W3], [ghajargar2022redhead_W7], [wan2024felt_W8], [liu2024personaflow_S2], [wang2024lave_M4], [qian2024shape_M5]
mixed-factorial design Latin Square [cai2023designaid_M1]
between-subjects design - [ege2024chatgpt_D10], [goodman2022lampost_W1], [kim2023cells_W9], [benharrak2024writer_W14]
Two Conditions - [xu2024jamplate_G3], [schuller2024generating_G9], [lee2024prompt_G18], [radensky2024scideator_S3], [banker2024machine_S4], [brade2023promptify_M3]
within-subjects design random counterbalancing [huang2024plantography_D14]
complete counterbalancing [paredes2024creative_G17], [reza2024abscribe_W13], [petridis2023anglekindling_W6], [pu2024ideasynth_S1], [wang2023popblends_M2]
mixed-factorial design complete counterbalancing [heyman2024supermind_G2]
between-subjects design - [chen2024asknaturenet_G10], [wang2023task_D8], [chen2024foundation_M8]
Three Conditions Latin Square [zhang2023visar_W4]
within-subjects design random counterbalancing [yuan2022wordcraft_W5]
between-subjects design - [goldi2024intelligent_W10]
Four Conditions within-subjects design - [wu2023iconshop_M7]

### v3

@@ 4.4 The Iteration of LLM-assisted Ideation Stages
Three studies [74, 93, 59] specifically involve selecting and processing individual ideas from idea clusters or multiple ideas. In these cases, the iterative process returns from the multi-idea evaluation and selection stage to the individual idea evaluation stage. This indicates that filtering and evaluating promising ideas at the idea cluster level is a promising approach.
Interestingly, two studies [67, 68] focus on the ideation of scientific research ideas. In these cases, idea iteration involves the exploration of new ideas and research starting points from multiple research ideas, representing a return to the scope definition stage of the ideation process. This finding suggests that in the context of scientific research, iterative ideation may involve revisiting and redefining the scope of the problem or research question, allowing for the exploration of novel research directions.
One study [35] employs the creation of different roles to facilitate product ideation. This approach represents an iteration that returns from the idea evaluation stage to the foundational materials collection stage, where the fundamental elements and perspectives that inform the ideation process are established. By iterating on the roles and personas involved in the ideation activity, this study highlights the importance of considering diverse viewpoints and contexts in the development of innovative product ideas.
@@ 5.1 Overview of Interaction Design in LLM-assisted Ideation Systems
[84] tool G2 Individual Text board & sidebar - Triggered Functions
@@
[83] tool G3 Individual Text board Add & Edit User-defined Prompt & Triggered Functions
[88] tool G4 Individual Text board - Triggered Functions
@@
[74] tool G5 Individual Text Canvas Add & Delete & Edit Triggered Functions
@@
[93] tool G7 Group Text Canvas & Chatbot Add & Delete & Edit Open Interaction & User-defined Prompt
@@
[91] tool G8 Group Text & Audio & Physicalisation Augmented Reality - Triggered Functions
[86] tool G10 Individual Text Canvas - Triggered Functions
@@
[87] tool G13 Individual Text Text Editor & Canvas Add & Edit User-defined Prompt & Triggered Functions
@@
[61] tool G14 Individual Text & Image board & sidebar - Triggered Functions
[90] tool G15 Group Text board - Triggered Functions
@@
[41] tool D7 Individual Text & Image Config Panel & Chatbot - Open Interaction
@@
[79] tool D9 Individual Text & Animation board & canvas - Automated Workflows
@@
[40] tool D13 Individual Text & Image Chatbot & Config Panel & Canvas Edit Open Interaction & User-defined Prompt
@@
[62] tool D14 Individual Text & Image Config Panel & Canvas Add & Delete & Edit Triggered Functions
@@
[56] tool W1 Individual Text Text Editor & Sidebar Add & Delete & Edit User-defined Prompt & Triggered Functions
@@
[43] tool W2 Individual Text Text Editor & Sidebar & Chatbot - Open Interaction & User-defined Prompt & Triggered Functions
@@
[50] tool W3 Individual Text & Sketch Text Editor & Canvas Add & Delete Triggered Functions
@@
[54] tool W4 Individual Text Text Editor & Canvas Add & Delete & Edit User-defined Prompt & Triggered Functions
@@
[47] tool W5 Individual Text Text Editor & Sidebar Add & Delete & Edit User-defined Prompt & Triggered Functions
[49] tool W6 Individual Text board & sidebar - Predefined Prompt
[51] tool W7 Individual Text Text Editor & Canvas - Triggered Functions
[48] tool W8 Individual Text Chatbot - Open Interaction
@@
[44] tool W9 Individual Text Config Panel & Text Editor Edit User-defined Prompt & Triggered Functions
[55] tool W10 Individual Text Text Editor & Sidebar - Triggered Functions
@@
[45] tool W13 Individual Text Text Editor & Sidebar Add & Delete & Edit User-defined Prompt & Triggered Functions
@@
[46] tool W14 Individual Text Text Editor & Sidebar Add & Delete & Edit User-defined Prompt & Triggered Functions
@@
[67] tool S1 Individual Text board & Canvas Add & Delete & Edit User-defined Prompt & Triggered Functions
@@
[68] tool S2 Individual Text Canvas Add & Delete & Edit User-defined Prompt & Triggered Functions
@@
[69] tool S3 Individual Text board Add & Delete User-defined Prompt & Triggered Functions
[57] tool M1 Individual Text & Image board - Triggered Functions
[58] tool M2 Individual Text & Image board - Predefined Prompt
@@
[59] tool M3 Individual Text & Image Config Panel & Canvas - User-defined Prompt & Triggered Functions
@@
[66] tool M4 Individual Text & Video Canvas & Chatbot - Open Interaction & User-defined Prompt & Triggered Functions
@@
[64] tool M5 Individual Text & Animation Config Panel & Canvas Edit Automated Workflows
[94] framework G6 Group Text & Audio - - Automated Workflows
[82] framework G12 Individual Text - - Automated Workflows
[85] framework G16 Individual Text - - Open Interaction & Predefined Prompt
[42] framework D5 Individual Text - - Automated Workflows
[80] framework D6 Group Text - - Automated Workflows
[72] framework S6 Individual Text - - Automated Workflows
[60] framework M8 Individual Text & Image - - Open Interaction
[95] method G1 Group Text - - Open Interaction
[89] method G9 Individual Text - - Open Interaction
[77] method G11 Individual Text - - Open Interaction
[92] method G17 Individual Audio & Physicalisation - - Open Interaction
[76] method G18 Individual Text - - Open Interaction
[35] method D1 Individual Text - - Open Interaction
[36] method D2 Group Text - - Open Interaction
[38] method D3 Individual Text - - Open Interaction
[39] method D4 Individual Text - - Open Interaction
[78] method D8 Individual Text - - Open Interaction & Predefined Prompt
[37] method D10 Group Text - - Open Interaction
[73] method D11 Group Text - - Open Interaction
[75] method D12 Individual Text & Image - - Open Interaction
[52] method W11 Individual Text - - Open Interaction
[53] method W12 Individual Text - - Open Interaction
[81] method W15 Individual Text - - Open Interaction
[70] method S4 Individual Text - - Open Interaction
[71] method S5 Individual Text - - Open Interaction
[63] method M6 Individual Text & Image - - Open Interaction
[65] method M7 Individual Text & Image - - Open Interaction
@@ 5.3 User Group
The prevalence of individually-oriented systems, such as LaMPost [56] and LAVE [66], demonstrates the emphasis on leveraging LLMs to enhance personal creativity and idea generation. These systems provide users with powerful tools to independently generate, refine, and evaluate ideas, augmenting their creative capabilities and enabling them to explore a wider range of possibilities.
However, the importance of group ideation cannot be overlooked, particularly in the context of large enterprises and organizations. Collaborative ideation offers several key benefits, including the exchange of diverse perspectives, the fostering of constructive criticism, and the cultivation of a shared sense of ownership over the generated ideas [93]. By bringing together individuals with different backgrounds, expertise, and viewpoints, team-based ideation can lead to more robust and innovative solutions.
Despite the critical role of group ideation, the interaction table reveals a notable scarcity of LLM-assisted tools specifically designed to support collaborative processes. Among the reviewed papers, only a few systems, such as the Collaborative Canvas [93] and the Conversational Swarm Intelligence system [94], are capable of supporting the collaborative processes in group ideation. This also indicates that developing collaborative systems for group ideation still faces unique challenges, which may include developing tools that support synchronous and asynchronous ideation, managing social dynamics, facilitating turn-taking, integrating the contributions of multiple participants, and ensuring equitable participation. There exist significant opportunities for future research and development. By focusing on the design and implementation of systems that cater to the specific needs of collaborative ideation, researchers can fully harness the potential of LLMs in supporting creative processes within teams and organizations.
@@ 5.4 Interaction Modality
The prevalence of text-based interaction in LLM-assisted ideation systems can be attributed to the initial development of LLMs, which focused heavily on text-based applications. This has led to the extensive use of LLMs in domains such as writing and other text-centric tasks. Many of the reviewed papers, such as those dealing with argumentative writing [54] and conceptual involvement [78], rely solely on text-based interaction to facilitate ideation processes.
However, in the realm of multimedia creation or domains involving multimedia elements, there is a growing trend towards combining text with various forms of media for interaction. For instance, image generation tasks [59] often employ a combination of text and image-based interaction, while video editing tasks [66] incorporate text and video modalities. Group ideation meetings [94] may leverage text and audio interaction to facilitate collaboration. Other interactive modalities explored in the reviewed papers include text and sketch [50], text and animation [79, 64], and even text and bodily movements [91].
Interestingly, only one paper in the reviewed set utilizes a non-text-based interaction modality. [92] explores the concept of assisting idea incubation in the context of commuter driving, employing a combination of driver body state and audio interaction. This unique approach highlights the potential for LLM-assisted ideation systems to extend beyond traditional text-based interactions and adapt to specific contextual requirements.
@@ 5.6 Idea Management Capability
Our analysis revealed that 17 tools (50%) developed idea components that considered users’ ability to manage ideas. Among these 17 tools, 10 incorporated all three functional components of adding, deleting, and editing ideas [74, 62, 45]. By offering a comprehensive set of idea management capabilities, these tools empower users to fully engage with the generated ideas, refining and shaping them to align with their creative goals.
Two tools [83, 87] focused on the addition and modification of ideas, while another two tools [50, 69] prioritized the addition and deletion of ideas. These variations in idea management capabilities reflect the different emphases and objectives of each tool, catering to specific aspects of the ideation process.
Interestingly, three tools [40, 44, 64] solely considered the modification of ideas. While the ability to edit ideas is undoubtedly valuable, the absence of addition and deletion functionalities may limit users’ capacity to fully explore the creative space and manage the generated ideas effectively.
@@ 5.7 Human-LLM Interaction Format
Predefined prompts are used in 4 out of 61 reviewed papers (7%). Two tool-type studies [49, 58] incorporate predefined prompts into their backend systems. The other two studies [85, 78] involve users directly interacting with LLMs using structured, predefined prompts.
@@ 6.1 User Study Design
One Condition - - [95], [83], [94], [77], [87], [61], [90], [85], [35], [80], [40], [43], [50], [51], [48], [68], [66], [64]
mixed-factorial design Latin Square [57]
between-subjects design - [37], [56], [44], [46]
Two Conditions - [83], [89], [76], [69], [70], [59]
within-subjects design random counterbalancing [62]
complete counterbalancing [92], [45], [49], [67], [58]
mixed-factorial design complete counterbalancing [84]
between-subjects design - [86], [78], [60]
Three Conditions Latin Square [54]
within-subjects design random counterbalancing [47]
between-subjects design - [55]
Four Conditions within-subjects design - [65]